a clinical report gave the following data from a blood analysis: iron 39 mg/dl, calcium 8.0 mg/dl cholesterol 224 mg/dl. how many grams of calcium are present in the blood of the patient if the blood volume is 2.6 liters? Please show all work thank you
Answer #1

Solution:-

Concentration of Ca = 8 mg/dl

First, we have to convert 8 mg/dl to g/L value.

1 mg = 10-3 g

1 dl = 10-1 L

Thus,

1 litre of solution contains = 0.08 g of Ca

So,

2.6 L of solution contains = (0.08 × 2.6) = 0.208 g of Ca

-------------------------------------------------------------